OpenStudy (sbuck98):

Someone help me answer these 4 problems, I will award a medal, and coins.

OpenStudy (sbuck98):

1.) Slope (m)= change in y over change in x. M= y2-41 over x2-x1 show your work on how you found the slop 2.) Find the equation using the point slope formula. Point slope formula is y-y1=m(x-x1 3.) .Give a example of a line parallel to your answer from question #2. 4.) Give a example of a line perpendicular to your answer from #2

OpenStudy (sbuck98):

@jim_thompson5910

OpenStudy (sbuck98):

@Chealyn98

OpenStudy (tkhunny):

Please show YOUR work. That's how we do things around here

OpenStudy (sbuck98):

I needed someone to HELP me do this.... That was my reason for posting asking FOR SOMEONE TO HELP ME..... I CAN'T SHOW WORK IF I HAVE NO IDEA HOW TO DO THIS.. @tkhunny

OpenStudy (sbuck98):

@TwilightSparkle2015 @TwelveBread49 @jackfrost33

OpenStudy (tkhunny):

1) It would help if you provided the entire problem. It appears that you need two points or you cannot solve the problems. 2) You CAN'T have NO idea if you are in the right class. You MUST have SOMETHING to show us. 3) Show YOUR work - every time.

OpenStudy (sbuck98):

Clearly I have no idea, I'm in the right class, and just started taking this class this week because I switched schools.... & I can't provide anymore information when that's all that was given to me...

OpenStudy (sbuck98):

@tkhunny

OpenStudy (tkhunny):

Okay, then you cannot solve the problem. The two points must be known. Since they are not known, you cannot proceed. Go get a better problem statement and we can talk. It is NOT clear that you have no idea. It is only clear that you have shown no work.